Logbooks Lab Maintenance Evaporator_1 Evaporator_2 Laser cutter Target Production Test-Stand RH-ISAC RH-Cyclotron RH-Meson Hall RH-Beamlines RH-ARIEL
  Evaporator_2, Page 16 of 35  Not logged in ELOG logo
ID Date Author Target Material Target Oven W.O# Source Status Subject Text Attachments
  389   Wednesday, September 27, 2017, 18:06 John Wongnp-SiC TARGET BOATDoneNS-npSiC & B3 pellets heat treatmentsWednesday, September 27, 2017, 18:05
 3 samples from NS-npSiC pellets,
pressured at 10T  (see lab book pg 53)
 Capture.JPG 
  388   Friday, September 15, 2017, 10:56 Laura LambertTa42218-10HP-SISDoneConditioning Ta#54 Friday, September 15, 2017, 09:44
Target consists of 470 discs.
 Capture.JPG 
  387   Tuesday, September 12, 2017, 13:08 Aaron Schmidt    Window shutter added Tuesday, September 12, 2017, 12:00:
The coated back window was changed
to a shuttered window.
  
  386   Monday, August 21, 2017, 17:32 John WongSiC TARGET BOATDoneCarbonization of sol-gel SiO2+CNTMonday, August 21, 2017, 17:09
Two pellets of sol-gel SiO2+CNT were
produced.  See lab book pg 44-46 for
  
  385   Tuesday, August 08, 2017, 15:00 Laura LambertEmpty  Problems resolvedWiring in PLC replaced re: IGP1 mis-communicating Tuesday, August 08, 2017, 14:57
See elog i.d's #384 and #381
Investigated wiring in PLC.
  
  384   Friday, August 04, 2017, 14:09 Laura Lambert 40282-5TARGET BOATDoneSintering La(OH)3/C pellets

    

       
            
  
  383   Friday, August 04, 2017, 10:52 Laura Lambert 40282-5TARGET BOATProblemsSintering La(OH)3/C pellets

    

       
            
 evap2_trip_8.4.2017.JPGevap2_trip_8.4.2017[2].JPGevap2_trip_8.4.2017[3].JPGevap2_trip_8.4.2017[4].JPG 
  382   Wednesday, August 02, 2017, 16:12 Marla Cervantes 40282-5 DoneSintering La(OH)3/C pellets Monday, July 31, 2017,  
Pellets previously pressed by
  
  381   Tuesday, August 01, 2017, 15:34 Laura Lambert   ProblemsIGP1 mis-communicating with EPICS - PLC problem?Monday, July 31, 2017, 15:24
Turned on BV and TP this afternoon.
Started ramping down, ~1min after
 evap2_trip_7.31.2017.JPG 
  380   Wednesday, July 19, 2017, 13:55 Marla Cervantes 40282-5TARGET BOATDoneSintering La(OH)3/C discs Wednesday, July 19, 2017, 13:44
Discs previously casted by Marla
and Pauline and weighed. 
 170724-LaC2-CarbothermalReduction-EPICS-PLOTS.PNGIMG-20170725-WA0018.jpg 
  379   Tuesday, July 18, 2017, 15:30 Laura Lambert   MaintenancePaddle wheel replacement for all evap2 water supplies Tuesday, July 18, 2017, 15:26
Problems occured with TGHT water
supply paddle wheel stopped and therefore
  
  378   Monday, July 10, 2017, 17:36 John Wongnp-SiC TARGET BOATProblemsWater Sensor Problem, TGHT Tripped. (np-SiC fibrous Pellet: Applied pressure vs Density)Monday, July 10, 2017, 17:28
Two np-SiC fibrous pellets with different
applied pressure:
 Graph3a.JPG 
  377   Tuesday, June 20, 2017, 16:07 Laura LambertTa42218-6HP-SISDoneConditioning Ta#53 Tuesday, June 20, 2017, 15:58
Target consists of 470 discs.
 Ta#53.JPG 
  376   Friday, June 16, 2017, 15:14 Laura LambertSiC40282-4TARGET BOATDoneSintering SiC pellets

    

       
            
 20170619_111816.jpg20170619_111848.jpg 
  375   Friday, June 16, 2017, 10:42 Marla CervantesSiC40282-4TARGET BOATDoneSintering SiC pellets

    

       
            
  
  374   Wednesday, June 14, 2017, 15:51 Laura LambertSiC40282-4TARGET BOATIn progressSintering SiC pellets Wednesday, June 14, 2017, 15:49
Pellets previously made by Marla
and weighed.
  
  373   Monday, June 05, 2017, 15:20 Laura LambertEmpty42218-15LP-SISDoneSintering TaC coating on UCx#21 container Monday, June 05, 2017, 15:20
Container coated with TaC using
  
  372   Tuesday, May 30, 2017, 14:40 Laura LambertTa42218-14LP-SISDoneConditioning Ta#52 Tuesday, May 30, 2017, 14:36
Target consists of 400 discs.
  
  371   Tuesday, May 16, 2017, 11:57 Laura LambertSiC41369-10HP-FEBIAD-CTLDoneSintering SiC#36 Tuesday, May 16, 2017, 10:56
Target container contains 460
 window_at_max_setpoint_after_1_day.jpg20170530_094047_Richtone(HDR).jpg20170530_094118_Richtone(HDR).jpg20170530_094137_Richtone(HDR).jpg 
  370   Thursday, May 11, 2017, 11:36 Laura LambertSiC39530-10HP-FEBIADDoneSintering SiC#36 - not SiC#36: now back-up [SiC#36 should be HP-FEBIAD-CTL!!!]

    

       
            
  
ELOG V2.9.2-2455